# Warm Inflation with Barrow Holographic Dark Energy

### v1

@@ Abstract
In this work we study the warm inflation mechanism in the presence of the Barrow holographic dark energy model. Warm inflation differs from other forms of inflation primarily in that it makes the assumption that radiation and inflaton exist and interact throughout the inflationary process. After the warming process, energy moves from the inflaton to the radiation as a result of the interaction, keeping the cosmos warm. Here we have set up the warm inflationary mechanism using Barrow holographic dark energy as the driving agent. Warm inflation has been explored in a high dissipative regime and interesting results have been obtained. It is seen that the Barrow holographic dark energy can successfully drive a warm inflationary scenario in the early universe. Finally, the model has been compared with the observational data and compliance has been found.
@@ 1 Introduction
Big-bang model is the most successful model in modern cosmology. According to the standard Big Bang cosmological model [1, 2, 3], the universe is either radiation or matter-dominated, which should lead to a decelerated expansion of the universe. Despite its success, the Big Bang theory may be incomplete in its classic form because it is not sufficient to solve some cosmological problems such as the flatness problem, horizon problem, and also the magnetic monopole problem [4, 5, 6, 7, 8]. These cosmological issues can be explained by the concept of inflation which was first proposed by A. Guth in 1981 [5]. Following inflation, the universe will enter a reheating phase, in which the inflaton decays into light particles, thermalizing the universe. By considering the cosmic expansion history from the time when the observed CMB scales escape the Hubble boundaries during inflation to when they re-enter it at a later time, it is conceivable to establish a link between the parameters of inflation and reheating. A review of inflationary cosmology can be found in [9]. Other notable developments on cosmological inflation can be found in [10, 11, 12, 13, 14, 15].
@@
The main idea of warm inflation [16] is that unlike cold inflation (the original inflation theory), there is an extra ”friction term” that acts as a regulator to fix the number density. In ordinary inflation, particle density exponentially reaches zero after inflation due to de-Sitter-type expansion. However, in warm inflation, we have specific characteristic temptation scales, such that the particle number does not go to zero due to coupling with another field. Warm inflation differs from other forms of inflation primarily in that it assumes that radiation and inflaton exist and interact throughout the inflationary process. After the warming process, energy moves from the inflaton to the radiation as a result of the interaction, keeping the cosmos warm. Consequently, warm inflation can lead to a very smooth phase of the universe, a radiation era, offering a novel solution to the graceful exit problem. Also, it should be noted that the cold inflation $\eta$ problem makes the flatness problem difficult to resolve. Warm inflation can naturally tackle the $\eta$ problem, hence offering much better reasoning for the flatness problem.
@@
It should also be noted that just like cold inflation affects the curvature perturbation, which in turn reciprocates to the CMB anisotropy, for warm inflation, similar calculations have been performed in [17]. This shows that during the viscous warm inflation time scale (typically 16 to 60 e-fold timing), one can work with some models motivated by field theory and string theory, which can indeed give the bound on the viscosity parameter (in adiabatic approximations) [18]. A study of warm inflation in both weak and strong dissipative regimes have been performed by Moss and Berera [19, 20] and as a consequence, one can show how phenomenological quantities such as scalar power spectrum behaves as a function of the dissipative parameter [21].
@@
Holographic dark energy is an alternative theory of dark energy, where we attempt to apply the holographic principle to the dark energy problem. Gerard ’t Hoof proposed the holographic principle [31, 32] inspired by the investigation of black hole thermodynamics [33]. The relationship between a quantum field theory’s greatest length and its ultraviolet cutoff [34] can result in holographic vacuum energy, which forms dark energy on cosmological scales [35, 36]. The holographic principle states that all of the information contained in a volume of a space can be portrayed as a hologram, that corresponds to a theory lying on the boundary of that space. The concept of holographic principle has been widely used in various fields of physics such as in nuclear physics to study the problems of quark-gluon plasma [37], in the field of condensed matter to study the problems of quark-gluon plasma [38], in the field of theoretical physics that lead to the idea of holographic entanglement entropy[39], in the field of cosmology to discuss the nature of de-Sitter space and inflation [40]. The holographic principle states that the universe’s horizon entropy is proportional to its area, comparable to the Bekenstein-Hawking entropy of a black hole. This is a key step in applying it to cosmology. Applying the holographic principle to the dark energy framework, a new model of dark energy known as the Holographic dark energy model (HDE) is formed. Very recently inspired by the illustrations of the Covid-19 virus, Barrow [41] showed that quantum-gravitational effects introduce the fractal features on the black-hole structure that leads to finite volume with infinite area. The corresponding black-hole entropy can be expressed as
@@ 2 Warm Inflationary mechanism
We start with the two fundamental equations of cosmology i.e. the Friedmann equations,
@@
where the subscripts ’r’ and ’in’ stands for radiation and the fluid that drives the inflation respectively. The conservation equation takes the form
$\dot{\rho_{in}}+3H\left(\rho_{in}+p_{in}\right)=-\Gamma\left(\rho_{in}+p_{in}\right)$ (2.3)
$\dot{\rho_{r}}+3H\left(\rho_{r}+p_{r}\right)=\Gamma\left(\rho_{r}+p_{r}\right)$ (2.4)
where ”r” stands for radiation and ”in” stands for the fluid that drives inflation. $\Gamma$ is referred to as the dissipation coefficient, and it may be constant, dependent on the scalar field or temperature $T_{r}$, or dependent on both the scalar field and temperature. The first slow-roll parameter is defined by,
@@
$\epsilon_{n+1}=\frac{\dot{\epsilon_{n}}}{H\epsilon}$ (2.6)
There are two types of inflationary models: ”warm Inflation” and ”cold Inflation”. In arm inflation, there is another type of slow roll parameter which is defined by
@@
It is known that both quantum and thermal fluctuations are present in case of warm inflation, and as long as $T>H$, the thermal fluctuations dominate. The two parameters, scalar spectral index ($n_{s}$) and tensor-to-scalar ratio ($r$) are widely used in inflationary scenarios. The scalar spectral index is defined by [48],
@@
and tensor-to-scalar can be expressed as
@@
where $\mathcal{P}_{t}$ denotes the amplitude of the tensor perturbation [49, 52].
Observational evidence indicates that the scalar spectral index should lie within the range $0.96$ to $0.9684$ and the upper limit of the parameter, the tensor-to-scalar ratio is $r<0.064$ [53]
@@ 3 Warm Inflation with Barrow Holographic dark energy
In this section, we consider that inflation is driven by holographic fluid. According to the holographic principle, holographic energy density is proportional to squared infrared cutoff $L_{IR}$. In the Barrow holographic dark energy model, the energy density is given by [47].
@@
where $C$ is the parameter whose dimension is given by $L^{-2-\Delta}$ and $R_{h}$ represents the future event horizen. If $\Delta=0$ then Eq.(3.1) takes the form $\rho_{BHDE}=CL^{-2}$ that provides the standard holographic dark energy model. Here $C=c^{2}{M_{p}}^{2}$ ($M_{p}$ is the Planck mass). The Granda-Oliveros (GO) cutoff is given by
$L^{-2}=\left(\alpha H^{2}+\beta\dot{H}\right)$ (3.2)
where $\alpha$ and $\beta$ are parameters. Implementing Eq.(3.2) in Eq.(3.1) we get energy density to be of the form
$\rho_{BHDE}=3{M_{p}}^{2}\left(\alpha H^{2}+\beta\dot{H}\right)^{1-\frac{\Delta}{2}}$ (3.3)
Since we have considered that inflation is driven by holographic fluid, therefore in our work $\rho_{inf}=\rho_{de}$. Therefore using Eq.(3.3) in Eq.(2.1) we get
$H^{2}=\frac{1}{3{M_{P}}^{2}}\left(\rho_{r}+3{M_{p}}^{2}(\alpha H^{2}+\beta\dot{H})^{1-\frac{\Delta}{2}}\right)$ (3.4)
Imposing quasi-stable production of radiation i.e $\dot{\rho}<<H\rho$ [54, 55] and using Eq.(2.4) and conservation equation we get
$4H\rho_{r}=\Gamma\left(\rho_{de}+p_{de}\right)$ (3.5)
@@
Inserting Eq.(3.6) in Eq.(3.4) $\dot{H}$ is obtained as
$\dot{H}=2\frac{1+Q}{Q}\left[\left(\alpha H^{2}+\beta\dot{H}\right)^{1-\frac{\Delta}{2}}-H^{2}\right]$ (3.7)
From Eq.(3.4) energy density for radiation is rewritten as
$\rho_{r}=3{M_{p}}^{2}\left[H^{2}-\left(\alpha H^{2}+\beta\dot{H}\right)^{1-\frac{\Delta}{2}}\right]$ (3.8)
@@
In particle physics models of inflation, the inflaton interacts with other fields rather than being isolated. Interactions may cause inflaton energy to dissipate into other degrees of freedom, resulting in a small percentage of the vacuum energy being converted to other forms of energy. Warm inflation involves a two-stage technique where dissipation produces particles with light degrees of freedom. In an expanding universe when relativistic particles thermalize fast enough, we can model their contribution as that of radiation:
$\rho_{r}=\sigma_{r}T^{4}$ (3.9)
@@
where $\sigma_{r}$ is the Stephen-Boltzman constant [56] and $T$ is the temperature of the radiation field. The Stephen-Boltzman constant can be expressed as $\sigma_{r}=\frac{{\pi}^{2}g}{30}$, where $g$ is the number of degrees of freedom of the radiation field.
Now the temperature can be obtained by equating Eq.(3.8) and Eq.(3.9)
$T^{4}=\frac{3M_{p}^{2}}{\sigma_{r}}\left[H^{2}-\left(\alpha H^{2}+\beta\dot{H}\right)^{1-\frac{\Delta}{2}}\right]$ (3.10)
@@
We know that the first slow roll parameter is defined as $\epsilon_{1}=-\frac{\dot{H}}{H^{2}}$. Therefore using Eq.(3.7) the first slow roll parameter can be expressed as
$\epsilon_{1}=-2\frac{1+Q}{Q}\left[H^{-\Delta}\left(\alpha+\beta\frac{\dot{H}}{H^{2}}\right)^{1-\frac{\Delta}{2}}-1\right]$ (3.11)
From Eq.(2.6) the second slow roll parameter is obtained as
@@
$\epsilon_{2}=\frac{\frac{(1+Q)}{Q}\left[(-\Delta)H^{-(\Delta+1)}\left(\alpha+\beta\frac{\dot{H}}{H^{2}}\right)^{1-\frac{\Delta}{2}}+H^{-\Delta}\left(1-\frac{\Delta}{2}\right)\left(\alpha+\beta\frac{\dot{H}}{H^{2}}\right)^{\frac{-\Delta}{2}}\left(\dot{\alpha}+\dot{\beta}\frac{\dot{H}}{H^{2}}+\beta\left(\frac{H\ddot{H}-2\dot{H}^{2}}{H^{3}}\right)\right)\right]-\frac{2\dot{Q}}{Q^{2}}}{\frac{1+Q}{Q}\left[H^{-\Delta}\left(\alpha+\beta\frac{\dot{H}}{H^{2}}\right)^{1-\frac{\Delta}{2}}-1\right]H}$ (3.12)
@@
Dissipative effects are important during the evolution of warm inflation. Friction causes the scalar field to dissipate into a thermal bath, resulting in dissipative effects. The dissipative coefficient $\Gamma$, a key quantity in supersymmetry, has been calculated from first principles in [57]. The co-efficient could be taken as constant but in a broader sense, it can be considered as a function of temperature $T$. Then, the power law form of the temperature can be considered as,
$\Gamma=B_{t}T^{m}$ (3.13)
where $B_{t}$ is constant. Therefore, using Eq.(3.10) $\Gamma$ can be reconstructed as
@@
$\Gamma=B_{t}\left(\frac{3{M_{p}}^{2}}{\sigma}\left[H^{2}-(\alpha H^{2}+\beta\dot{H})^{1-\frac{\Delta}{2}}\right]\right)^{\frac{m}{4}}$ (3.14)

### v2

@@ Abstract
In this work we study the warm inflation mechanism in the presence of the Barrow holographic dark energy model. Warm inflation differs from other forms of inflation primarily in that it assumes that radiation and inflaton fields exist and interact throughout the inflationary process. After the warming process, energy moves from the inflaton to the radiation as a result of the interaction, keeping the cosmos warm. Here we have set up the warm inflationary mechanism using Barrow holographic dark energy as the driving agent. Warm inflation has been explored in a highly dissipative regime and interesting results have been obtained. It is seen that the Barrow holographic dark energy can successfully drive a warm inflationary scenario in the early universe. Finally, the model was compared with the observational data, and compliance was found.
@@ 1 Introduction
The Big Bang model is the most successful model in modern cosmology. Despite its success, the Big Bang theory [1, 2, 3] may be incomplete in its classic form because it is not sufficient to solve some cosmological problems such as the flatness problem, horizon problem, the magnetic monopole problem [4, 5, 6, 7, 8] and the accelerated expansion of the universe [9, 10]. The flatness, horizon, and magnetic monopole problem can be explained by the concept of inflation which was first proposed by A. Guth in 1981 [5]. Following inflation, the universe will enter a reheating phase, in which the inflaton decays into light particles, thermalizing the universe. By considering the cosmic expansion history from the time when the observed CMB scales escape the Hubble boundaries during inflation to when they re-enter it at a later time, it is conceivable to establish a link between the parameters of inflation and reheating. A review of inflationary cosmology can be found in [11]. Other notable developments on cosmological inflation can be found in [12, 13, 14, 15, 16, 17, 18, 19].
@@
The main idea of warm inflation [20] is that unlike cold inflation (the original inflation theory), there is an extra ”friction term” that acts as a regulator to fix the number density. In ordinary inflation, particle density exponentially reaches zero after inflation due to de-Sitter-type expansion. However, in warm inflation, we have specific characteristic temptation scales, such that the particle number does not go to zero due to coupling with another field. Warm inflation differs from other forms of inflation primarily in that it assumes that radiation and inflaton exist and interact throughout the inflationary process. After the warming process, energy moves from the inflaton to the radiation as a result of the interaction, keeping the cosmos warm. Consequently, warm inflation can lead to a very smooth phase of the universe, a radiation era, offering a novel solution to the graceful exit problem. One of the biggest challenges to developing a workable inflationary model is the $\eta$ problem [21, 22, 23]. Long-term slow-roll inflation is impossible due to Planck mass-suppressed adjustments to the inflaton potential, which typically result in inflaton masses of the order of the Hubble scale and yield contributions of order unity to the $\eta$ slow roll parameter. It should be noted that the cold inflation $\eta$ problem makes the flatness problem difficult to resolve. Warm inflation can naturally tackle the $\eta$ problem, hence offering much better reasoning for the flatness problem [24]. This is a great motivation for studying warm inflation.
@@
It should also be noted that just like cold inflation affects the curvature perturbation, which in turn reciprocates to the CMB anisotropy, for warm inflation, similar calculations have been performed in [25]. This shows that during the warm inflation time scale (typically 16 to 60 e-fold timing), one can work with some models motivated by field theory and string theory, which can indeed give the bound on the dissipation parameter (in adiabatic approximations) [26]. A study of warm inflation in both weak and strong dissipative regimes have been performed by Moss and Berera [27, 28] and as a consequence, one can show how phenomenological quantities such as scalar power spectrum behaves as a function of the dissipative parameter [29].
@@
Holographic dark energy is an alternative theory of dark energy, where we attempt to apply the holographic principle to the dark energy problem. Gerard ’t Hooft proposed the holographic principle [39, 40] inspired by the investigation of black hole thermodynamics [41]. The relationship between a quantum field theory’s greatest length and its ultraviolet cutoff [42] can result in holographic vacuum energy, which forms dark energy on cosmological scales [43, 44]. The holographic principle states that all of the information contained in a volume of a space can be portrayed as a hologram, that corresponds to a theory lying on the boundary of that space. The concept of holographic principle has been widely used in various fields of physics such as in nuclear physics to study the problems of quark-gluon plasma [45], in the field of condensed matter to study the problems of strongly correlated systems [46], in the field of theoretical physics that lead to the idea of holographic entanglement entropy[47], in the field of cosmology to discuss the nature of de-Sitter space and inflation [48]. The holographic principle states that the universe’s horizon entropy is proportional to its area, comparable to the Bekenstein-Hawking entropy of a black hole. This is a key step in applying it to cosmology. Applying the holographic principle to the dark energy framework, a new model of dark energy known as the Holographic dark energy model (HDE) is formed [49, 50]. In recent years Holographic dark energy has been a very lucrative candidate for explaining late-time cosmic acceleration. Rooted in the Holographic principle from string theory holographic dark energy can explain late-time cosmic acceleration and is very consistent with the observational data [51, 52, 53, 54, 55, 56]. The current landscape of holographic dark energy research is not just bound in general relativity, as we know that in a high-energy regime, Einstein’s gravity is incapable of being consistent with quantum gravity. As a consequence, various authors have studied holographic dark energy in the context of various modified gravity such as Brans-Dicke gravity [57], $f(R,G)$ gravity [58], $f(Q)$ gravity [59], $f(T,T_{G})$ gravity [60], $f(R,T)$ gravity [61], $f(Q,T)$ gravity [62], $f(Q,C)$ gravity [63] etc. Very recently inspired by the illustrations of the Covid-19 virus, Barrow [64] showed that quantum-gravitational effects introduce the fractal features on the black-hole structure that leads to finite volume with infinite area. The corresponding black-hole entropy can be expressed as
@@ 2 Warm Inflationary mechanism
To formulate the warm inflationary dynamics mathematically, we begin with the Einstein field equations in a spatially flat Friedmann-Robertson-Walker (FRW) universe. The metric is given by
$ds^{2}=dt^{2}-a^{2}(t)(dx^{2}+dy^{2}+dz^{2})$ (2.1)
where $a(t)$ is the scale factor. In this framework, the total energy density comprises contributions from both the inflaton field and the radiation fluid. The evolution of this system is governed by the two fundamental equations of cosmology i.e. the Friedmann equations,
@@
where the subscripts ’$r$’ and ’$in$’ stands for radiation and the fluid that drives the inflation respectively. The conservation equation takes the form,
$\dot{\rho_{in}}+3H\left(\rho_{in}+p_{in}\right)=-\Gamma\left(\rho_{in}+p_{in}\right),$ (2.4)
$\dot{\rho_{r}}+3H\left(\rho_{r}+p_{r}\right)=\Gamma\left(\rho_{in}+p_{in}\right),$ (2.5)
where $\Gamma$ is referred to as the dissipation coefficient, and it may be constant, dependent on the scalar field or temperature $T_{r}$, or dependent on both the scalar field and temperature. The first slow-roll parameter is defined by,
@@
$\epsilon_{n+1}=\frac{\dot{\epsilon_{n}}}{H\epsilon_{n}}$ (2.7)
There are two types of inflationary models: ”warm Inflation” and ”cold Inflation”. In warm inflation, there is another type of slow roll parameter which is defined by
@@
It is known that both quantum and thermal fluctuations are present in case of warm inflation, and as long as $T>H$, the thermal fluctuations dominate. The function $G(Q)$ is the growth of inflation fluctuations due to the coupling to radiation and can be determined numerically [36]. By extending the study of [76] we numerically obtain Eq.(2.11). The two parameters, scalar spectral index ($n_{s}$) and tensor-to-scalar ratio ($r$) are widely used in inflationary scenarios. The scalar spectral index is defined by [74],
@@
where $k=aH$. We can write $d~{}ln~{}k=\frac{dlnk}{dN}dN$. Therefore $\frac{dlnk}{dN}$ can be expressed as dissipative parameter $Q$ and first slow roll parameter $\epsilon_{1}$ i.e $\frac{dlnk}{dN}=1-\frac{\epsilon_{1}}{1+Q}$ [70, 83]. When $Q\to 0$ and $T\to 0$ the spectral index for cold inflation reduces to $n_{s}=1-6\epsilon+2\eta$ (for detailed calculation see [78]).
During inflationary expansion, both scalar (density) perturbations and tensor (gravitational wave) perturbations are generated. These primordial perturbations are crucial as they seed the large-scale structure formation in the universe and leave distinctive imprints on the cosmic microwave background radiation. An important observational parameter that characterizes these perturbations is the tensor-to-scalar ratio, which measures the relative strength of tensor-to-scalar perturbations. This ratio is defined as,
@@
where $\mathcal{P}_{t}$ denotes the amplitude of the tensor perturbation which is defined by $\mathcal{P}_{t}=\frac{2H^{2}}{{\pi}^{2}{M_{p}}^{2}}$ [36, 77].
Observational evidence indicates that the scalar spectral index should lie within the range $0.96$ to $0.9684$ and the upper limit of the parameter, the tensor-to-scalar ratio is $r<0.064$ [79].
@@ 3 Warm Inflation with Barrow Holographic dark energy
In this section, we consider that inflation is driven by a holographic fluid. According to the holographic principle, holographic energy density is proportional to squared infrared cutoff $L_{IR}$. In the Barrow holographic dark energy model, the energy density is given by [73].
@@
where $C$ is the parameter whose dimension is given by $L^{-2-\Delta}$ If $\Delta=0$ then Eq.(3.1) takes the form $\rho_{BHDE}=CL^{-2}$ that provides the standard holographic dark energy model. Here $C=3c^{2}{M_{p}}^{2}$ ($M_{p}$ is the Planck mass). The Granda-Oliveros (GO) cutoff is given by [71, 72],
$L^{-2}=\left(\alpha H^{2}+\beta\dot{H}\right),$ (3.2)
where $\alpha$ and $\beta$ are parameters. Implementing Eq.(3.2) in Eq.(3.1) we get energy density to be of the form [72],
$\rho_{BHDE}=3{M_{p}}^{2}\left(\alpha H^{2}+\beta\dot{H}\right)^{1-\frac{\Delta}{2}}.$ (3.3)
Here $C$ is replaced by $3{M_{p}}^{2}$ (considering c=1) [72]. Since we have considered that inflation is driven by holographic fluid, therefore in our work $\rho_{in}=\rho_{de}$. Therefore using Eq.(3.3) in Eq.(2.2) we get
$H^{2}=\frac{1}{3{M_{P}}^{2}}\left(\rho_{r}+3{M_{p}}^{2}(\alpha H^{2}+\beta\dot{H})^{1-\frac{\Delta}{2}}\right).$ (3.4)
Imposing quasi-stable production of radiation i.e $\dot{\rho}_{r}<<H\rho_{r}$ [80, 81] and using Eq.(2.5) and conservation equation we get
$4H\rho_{r}=\Gamma\left(\rho_{de}+p_{de}\right).$ (3.5)
@@
Inserting Eq.(3.6) in Eq.(3.4) $\dot{H}$ is obtained as,
$\dot{H}=2\frac{1+Q}{Q}\left[\left(\alpha H^{2}+\beta\dot{H}\right)^{1-\frac{\Delta}{2}}-H^{2}\right].$ (3.7)
From Eq.(3.4), energy density for radiation is rewritten as,
$\rho_{r}=3{M_{p}}^{2}\left[H^{2}-\left(\alpha H^{2}+\beta\dot{H}\right)^{1-\frac{\Delta}{2}}\right].$ (3.8)
@@
In particle physics models of inflation, the inflaton interacts with other fields rather than being isolated. Interactions may cause inflaton energy to dissipate into other degrees of freedom, resulting in a small percentage of the vacuum energy being converted to other forms of energy. Warm inflation involves a two-stage technique where dissipation produces particles with light degrees of freedom. In an expanding universe when relativistic particles thermalize fast enough, we can model their contribution as that of radiation,
$\rho_{r}=\sigma_{r}T^{4},$ (3.9)
@@
where $\sigma_{r}$ is the Stephen-Boltzmann constant [34] and $T$ is the temperature of the radiation field. The Stephen-Boltzmann constant can be expressed as $\sigma_{r}=\frac{{\pi}^{2}g}{30}$, where $g$ is the number of degrees of freedom of the radiation field.
Now the temperature can be obtained by equating Eq.(3.8) and Eq.(3.9),
$T^{4}=\frac{3M_{p}^{2}}{\sigma_{r}}\left[H^{2}-\left(\alpha H^{2}+\beta\dot{H}\right)^{1-\frac{\Delta}{2}}\right].$ (3.10)
@@
We know that the first slow roll parameter is defined as $\epsilon_{1}=-\frac{\dot{H}}{H^{2}}$. Therefore using Eq.(3.7) the first slow roll parameter can be expressed as,
$\epsilon_{1}=-2\frac{1+Q}{Q}\left[H^{-\Delta}\left(\alpha+\beta\frac{\dot{H}}{H^{2}}\right)^{1-\frac{\Delta}{2}}-1\right].$ (3.11)
From Eq.(2.7) the second slow roll parameter is obtained as,
@@
$\epsilon_{2}=\frac{\frac{(1+Q)}{Q}\left[(-\Delta)H^{-(\Delta+1)}\left(\alpha+\beta\frac{\dot{H}}{H^{2}}\right)^{1-\frac{\Delta}{2}}+H^{-\Delta}\left(1-\frac{\Delta}{2}\right)\left(\alpha+\beta\frac{\dot{H}}{H^{2}}\right)^{\frac{-\Delta}{2}}\left(\dot{\alpha}+\dot{\beta}\frac{\dot{H}}{H^{2}}+\beta\left(\frac{H\ddot{H}-2\dot{H}^{2}}{H^{3}}\right)\right)\right]-\frac{2\dot{Q}}{Q^{2}}}{\frac{1+Q}{Q}\left[H^{-\Delta}\left(\alpha+\beta\frac{\dot{H}}{H^{2}}\right)^{1-\frac{\Delta}{2}}-1\right]H}.$ (3.12)
@@
Dissipative effects are important during the evolution of warm inflation. Friction causes the scalar field to dissipate into a thermal bath, resulting in dissipative effects. The dissipative coefficient $\Gamma$, has been calculated from the first principles in [82, 31] in the context of supersymmetry. However, there are other ways to construct warm inflation which does not require the use of supersymmetry as done in [83]. The co-efficient could be taken as constant but in a broader sense, it can be considered as a function of temperature $T$. Then, the power law form of the temperature can be considered as,
$\Gamma=B_{t}T^{m},$ (3.13)
where $B_{t}$ is a constant. Therefore, using Eq.(3.10) $\Gamma$ can be reconstructed as,
@@
$\Gamma=B_{t}\left[\frac{3{M_{p}}^{2}}{\sigma_{r}}\left(H^{2}-(\alpha H^{2}+\beta\dot{H})^{1-\frac{\Delta}{2}}\right)\right]^{\frac{m}{4}}.$ (3.14)